4(2x-1)-1=6. how do you find out what x is? please help me!
Answer by ikdeep(226)

4(2x-1)-1=6
8x - 4 - 1 = 6
8x - 5 = 6
8x = 6 + 5
8x = 11
x = 11/8
To verify your answer you can put the value of x in the given equation and you will get LHS=RHS.
Please feel free to revert back for any further queries.